OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined. frac{t^3 -9t}{t^2 -36}

OpenStudy (anonymous):

\[\frac {t^3 -9t}{t^2 -36}\]

OpenStudy (anonymous):

Where t^2 - 36 = 0 (t+6)(t-6) = 0 Where t = -6, 6

OpenStudy (chaise):

Well, actually. The denominator cannot equal 0 otherwise the equatioon becommes undefined.

OpenStudy (chaise):

I don't even know what the question is asking. Can someone explain so I can help?

OpenStudy (anonymous):

The question is asking "The numbers for which the rational expression is undefined are _

OpenStudy (chaise):

So to make the expression undefined you would need the denominator to equal zero. That person was right, lol, my bad.

OpenStudy (anonymous):

It's cool :) Thanks for the reassurance

OpenStudy (anonymous):

\[(t^3−3)/(t^2-36) =t (t−3)(t+3) /(t-6)(t+6) \]